The mean height of a random sample of 133 people is 155 cm and a standard deviation of 15 cm. The proportion of persons taller than 200 cm


1
Expert's answer
2022-05-03T14:30:42-0400
"P(X>200)=1-P(Z\le\dfrac{200-155}{15/\sqrt{133}})"

"\approx0"

Hence, from the population of 133 people, No one is taller than 200 cm.